Independent random samples were selected from two quantitative populations, with sample sizes, means, and standard deviations given below.

n1 = n2 = 60

x1 = 125.3

x2 = 123.4

s1 = 5.7

s2 = 6.1

a) Construct a 95% confidence interval for the difference in the population means (μ1 − μ2). (Round your answers to two decimal places.) to

b) Find a point estimate for the difference in the population means.

c) Calculate the margin of error. (Round your answer to two decimal places.)
Answer #1

Sample Mean 1 (X1) = 125.3 Sample Standard Deviation 1 (81) = 5.7 Sample Size 1 (N1) = 60 Sample Mean 2 (X2) = 123.4 Sample S

2.

Point Estimate = 125.3 - 123.4 = 1.9

3.

Margin of error = tc * std error

Margin of error = 1.98 * 1.078 = 2.13

Dear student,
I am waiting for your feedback. I have given my 100% to solve your queries. If you satisfied with my answer then please please like this.
Thank You